please help me solve

2b^ - 4ac = bc where a = -2, b = 4, c = 3

Answer by checkley71(8403)

2B^(2?)-4AC=BC
2*4^2-4*-2*3=4*3
2*16+24=12
32+24=12
56=12 SOMETHING IS WRONG HERE!!!!!!!
2B-4AC=BC
2*4-4*-2*3=4*3
8+24=12
32=12 ALSO DOESN'T MATCH.